A nursery wants to mix two types of seeds, one type sells for 12 dollars per pound and the other sells at 20 per pound. To obtain 100 pounds of a mixture at 14 dollars per pound, how many pounds of each seeds are needed?
Answer by stanbon(75887)

A nursery wants to mix two types of seeds, one type sells for 12 dollars per pound and the other sells at 20 per pound. To obtain 100 pounds of a mixture at 14 dollars per pound, how many pounds of each seeds are needed?
-----------------
Let amont of $12 seed be x lbs; value is 12x dollars
-------------------------
Amount of $20 seed is "100-x" ; value is 20(100-x) = (2000-20x) dollars
------------------------
Amt of mixture = 100 lbs; value = 14*100 = 1400 dollars
-----------------------
EQUATION:
value + value = value
12x + 2000-20x = 1400
-8x = -600
x = 75 lbs (amount of $12 see in the mixture)
100-x = 25 lbs (amount of $20 see in the mixture)
===============
Cheers,
Stan H.